Run Coordination Meetings: Spring 2017 Run

From GlueXWiki
Jump to: navigation, search

Meeting General Agenda

Run Planning Meeting Notes, Mar 9 - Mar 15, 2017

Run Planning Meeting Notes, Mar 2 - Mar 8, 2017

Run Planning Meeting Notes, Feb 23- Mar 1, 2017

Run Planning Meeting Notes, Feb 16-Feb 22, 2017

Run Planning Meeting Notes, Feb 9-Feb 15, 2017

Connect Information

To connect from the outside by phone please do the following

1.) To join via a Web Browser, go to the page [1] https://bluejeans.com/660743227.

2.) To join via Polycom room system go to the IP Address: 199.48.152.152 (bjn.vc) and enter the meeting ID: 660743227.

3.) To join via phone, use one of the following numbers and the Conference ID: 660743227.

4.) The moderator code is 8394

Dial Either +1 408 740 7256 or +1 888 240 2560 for US or Canada

Specific instructions for connecting from JLab's Hall D conference room:

  • Turn polycon on if necessary (do it before turning the computer on)
  • With the polycon, place a call at 199.48.152.152
  • Press # to enable the polycom keypad, then enter the meeting id: 660743227 and #
  • You may have to unmute the microphone: #*4
  • Turn the computer on if needed

Beam and Hall configuration

  • Beam energy expected: 11.640 GeV (value from Fall 2017 measurement).
  • Solenoid on at 1350A. Any field changes should be avoided if possible.
  • CW Beam current:
    • Range: 1 nA-1.5 μA. 250 MHz frequency.
    • Standard production running expected at ~200 nA on 4·10-4 R.L. radiator.
    • RSAD limit for most of the run, excluding high intensity tests: 250 nA on 4·10-4 R.L. radiator.
  • FFB on or Position Slow Lock on (FFB doesn't operate below 50nA).
  • Radiators:
    • Goniometer diamond radiators:
      • J1a50       (50 μm - 4.1·10-4 R.L., 4×4 mm²) - old crystal - excellent rocking curve
      • JD70-100 (58 μm - 4.8·10-4 R.L., 7×7 mm²) - new crystal - excellent rocking curve
      • JD70-105 (47 μm - 3.9·10-4 R.L., 7×7 mm²) - new crystal - good rocking curve
      • JD70-104 (17 μm - 1.4·10-4 R.L., 7×7 mm²) - new crystal - wider rocking curve - marginally usable
    • We also have one 10 μm Al radiator.
    • Amorphous
      • 1.5 µm Al (1.7·10-5 R.L.)
      • 10 µm Al (11.2·10-5 R.L.)
      • 30 µm Al (33.7·10-5 R.L.)
  • Tagger quadrupole on (-4.2 A).
  • Collimator hole: Both 5 mm diameter and 3.4 mm diameter. Default running: 5.0 mm

General Information

This document describes the run plan for the Hall D Spring 2017 run.

  • Schedule for the Spring 2017 run:
  1. Jan. 26th-Jan. 29th: Electron beam restoration
  2. Jan. 30th-Mar. 22nd: Hall D Spring run.
  3. Every Tuesdays: 12h of beam study. Every Thursdays, if necessary: 8h of RF recovery.
  • overall plan and priorities:
  1. The first priority is to do the Halls B and C KPP (key performance parameters).
  2. The second priority is to finish commissioning the 5-pass separator and to ensure we can reliably run with it. This is in preparation for the Fall 17 run when all halls will be a 5-pass (Hall D at 5.5 pass).
    • The above implies that Hall D should be run continuously at 5.5 pass and hence, the Hall D physics run is tie to priority #2.
    • Hall B starts its KPP run (3-pass) on Feb. 3rd. With smooth running, it should achieve the KPP goals in a day or two. Then, it shuts down.
    • Hall C starts its KPP run (3-pass) around March 7th. It will then continue running at 1-pass for spectrometer commissioning (if its magnets are cold and ready). Then, if time permit before the end of the spring run, they may start their physics program at 5-pass.
    • Hall A will be used for its beam-dump when Hall C is down, so that a high current beam can be run in the machine when Hall B and D are running. This high current beam will stabilize the beam orbit in the machine. Hall A may take advantage of receiving beam by opportunistically running its argon experiment. In that case, it will be at 1-pass.
  • Experiment: GlueX E12-06-102 Production
  • Planned tests:
    • High-intensity GlueX test (invasive, 3 shifts)
    • TPol low energy photon run (invasive, 2 shifts)
    • ComCAL(PrimeX) prototype test (parasitic)
    • TRD test (parasitic)
  • Expected Staffing and responsibilities:
    • The Run Coordinator oversees the commissioning:
      • Thur. Jan. 26 - Sun. Jan. 29 (4 days): A. Deur (accelerator restoration)
      • Mon. Jan. 30 - Wed. Feb. 8 (10 days): A. Deur
      • Wed. Feb. 8 - Wed. Feb. 15 (7 days): M. Dalton (with Mon. 6th - Wed. 8th overlap)
      • Wed. Feb. 15 - Wed. Feb. 22 (7 days): A. Ostrovidov (with Mon. 13th - Wed. 15th overlap)
      • Wed. Feb. 22 - Wed. Mar. 1 (7 days): W. Boeglin (with Mon. 20th - Wed. 22nd overlap)
      • Wed. Mar. 1 - Wed. Mar. 8: (7 days): S. Dobbs (with Mon. 27th - Wed. 1st overlap)
      • Wed. Mar. 8 - Wed. Mar. 15: (7 days): R. Jones (with Mon. 6th - Wed. 8th overlap)
      • Wed. Mar. 15 - Wed. Mar. 22: (7 days): D. Lawrence (with Mon. 13th - Wed. 15th overlap)


Runplan

  1. Verify electron beam quality and establish photon beam (assume 1 day).
    • Ion Chambers trip threshold must be set using Al. radiators
    • Tune electron beam with radiator retracted
      • Radiator (amorphous and diamond) must be retracted. Collimator should be in blocking position.
        • Typical levels are (for radiator retracted, collimator fully blocking, electron beam current 50 nA)
          • RAD102_P1 ~ 10 mrad/hr (tagger area, gammas, between tagger and dump)
          • RAD102_P2 ~ 30 mrad/hr (tagger area, gammas, on goniometer).
          • RAD102_P3 ~ 0.1 mrem/hr (tagger area, neutrons, near electronics racks)
          • RAD508_P1 < 0.1 mrad/hr (collimator cave, gammas).
          • RAD508_P2 ~ 0. mrad/hr (Hall. Neutrons).
      • Hall D tagger CARMS, Ion Chamber and BLM threshold for beam trips should not be changed without approval of Hall D leader.
    • Establish good photon beam
    • Make sure tagger quadrupole is on (-4.2 A).
      • Insert 2*10-5 RL radiator. (Call MCC and ask to mask the FSD each time the radiator is moved)
        • Typical levels are (for 2*10-5 radiator, electron beam current 50 nA)
          • RAD102_P1= 15 mrad/hr (tagger area, gammas, between tagger and dump)
          • RAD102_P2= 30 mrad/hr (tagger area, gammas, on goniometer)
          • RAD102_P3= ~2 mrem/hr (tagger area, neutrons, near electronics racks)
          • RAD508_P1= ~10 mrad/hr (collimator cave, gammas).
          • RAD508_P2 < ~ 0. mrad/hr (Hall. Neutrons).
    • Once radiation levels are good, proceed with harp scans (Hovanes/Todd. 15 min)
      • Take tagger harp scans. Check consistency with Accelerator's harps. Study beam focusing
    • Radiation level study (1h):
      • Run 5 min without radiator (Call MCC and ask to mask the FSD each time the radiator is moved). Beam current: 50nA. Repeat at 100 nA and 200 nA
      • Run 5 min with the 10-4 radiator (Call MCC). Beam current: 50nA. Repeat at 100 nA and 200 nA
      • Run 3 min with the 3*10-4 radiator (Call MCC). Beam current: 50nA. Repeat at 100 nA and 200 nA
    • Active Collimator calibration. (20 + 20 mins). Call Hovanes.
      • re-Insert the 10-4 radiator (Call MCC). Use about 100 nA beam current. This will provide the Act. Col. nominal positions for its high gain values.
      • Determine maximum photon beam transmission by doing 2D scans using the collimator x/y motions (Hovanes. 1h)
      • Repeat procedure with 2*10-5 radiator (Call MCC) and 50 nA beam current. This will provide the Act. Col. nominal positions for its low gain values.
      • Record on the white board what is the beam position on 5C11b and Active Collimator for optimal photon transmission, both for low and high Act. Col. gain.
      • Retract profiler.
    • Verify with MCC that position slow lock is back on (locking on Act. Col. and for values corresponding to optimal photon flux).
  2. Detector and beamline checkout.
    • Reestablish L1 trigger Alex. Somov. 3h).
    • Reestablish DAQ (Sergei, 2h).
    • Microscope commissioning. Alex Barnes (2h).
    • General detector calibration (Sean Dobbs . 4h). Use 10-4 radiator. Call Sean when getting ready for this task (at any hour): (703) 887-1776.
      • Run 15 min in Long Mode.
    • From here on, the tasks below are interspersed with data production
  3. Realign 60 μm diamond (JD70-100). Later on, align new 20 μm diamond JD70-104 (Hovanes, Paul Mattione. 12h/diamond).
  4. L1 trigger test for high-rate performance tests (Alex Somov. 4h+4h+4h+4h. Use up to 1.5 μA. DC off if necessary)
  5. Run PS at lower B-field (450A) for the TPol can check the low photon energy tail polarization. Use specific low-energy polarimetry trigger. (Alex Somov/Nathan Sparks. (12h+12h)
  6. DAQ high-rate performance tests (Sergei Furletov 2h)
  7. L3 trigger test (David Lawrence. 4h)
  8. Straight track run (16h, nominally at the end of the Spring run)
    • Ramp down solenoid magnet (2h)
    • Take straight track run data (Lubomir)
  9. Gluex data production
    • Log in run information in the shift summary (make a run list). For each production run, log useful comments on DAQ comment window, see [2]
    • Tagger quadrupole on
    • Harp scans once a day: 15 min.
    • Empty target run every week (2h: 1- h to empty/fill back the target, 1h running). Standard production current.
    • Long mode 10 min run. Once a week (20 min).
    • Amorphous run. Every day. 1h Standard production current. Time to switch from pol. to unpol.: 10min. time to switch from unpol. to pol.: 30min.
    • TAC runs. 2 TAC runs (one at the beginning, one in the middle of the run). (Hovanes, Alex Somov, 10h: 8h run with thick converter. 1- LH2 target empty). Use 3.4mm collimator with 1mm plug in (if available).
    • Physics production data with diamonds and 5 mm hole. Ratio of para./transverse/45o/135o data:25/25/25/25. 2h runs and no more, as it may be hard to correct for calibration drifts. Switch polarization every run. Time to switch: 10min. DAQ start run overhead: 3min.
      • Start run with 60 μm diamond (JD70-100). When new 20 μm diamonds is aligned, do 2 evenings of production to assess its quality and decide what diamond should be used.
      • Luminosity:
        • Run at about 200 nA (for 60 μm diamond JD70-100).
  1. Parasitic run for PrimEx detector at the end during B=0 straight track run.
  2. Parasitic TRD run

Hall D white board (current run plan)

Hall D counting room white board

Runplan time charts


Accelerator status

Hot Checkout

  1. Hot Check OUT Status 100% ready ; 0% checked ; 0% not ready

Readiness for Fall 16 run

  • Solenoid (Eugene/Beni)
    • Problem with ground fault detector is fixed.
    • Solenoid to be ramp-up on Monday morning.
  • Surveys (Tim/Tom)
  • DAQ (Sergei)
    • Found the reason for the corrupted files problem and fixed it. We can run in the same way as Fall 16, or install a hardware controller which redundancy would make the data taking safer.
    • New version of CODA is available.
  • Trigger (Alex S.)
  • Beam line
    • Diamond status: They are ready
    • Tagger hodoscope (Nathan/Franz)
      • Took some data. Checks are good so far.
    • Tagger microscope (Alex B./Richard)
      • New fibers are now installed. Need to check for dead channels.
    • Goniometer (Hovanes)
      • Need to reglue the diamond that fell off (J1A50 50 μm, 4×4 mm² diamond. It is unclear when it fell. Maybe this is related to the fact that Hovanes and Paul did not find the diamond with the e- beam)
    • Collimator, Active collimator, profiler, flux monitor (Alex B./Hovanes).
      • Flux monitor not checked yet. Working on profiler software to be able to mask noisy channels.
    • PS system (Alex S.)
    • Polarimeter (M. Dugger/Nathan)
      • Turned it on. Need to check that the data are good.
    • Target (Tim/Alexandre D.)
      • Will be checked by the target group today (Thursday 26th).
    • TAC (Hovanes)
      • Not checked yet.
    • Slow control (Hovanes)
      • Not checked yet.
    • RF (Paul Mattione)
    • Radiation monitors
      • OSL and dosimetry rods installed around goniometer. Photon probe put back there too.
      • Downstream photon probe is broken.
      • Radiation monitors are ready and checked.
  • Detector systems
    • ST (Mark Ito)
      • ST has been turned on. Taking data for HCO.
    • FDC (Lubomir)
      • Need to start to get cosmic data
    • CDC (Beni)
      • There may be a leak in the gas system There was no leak.
    • BCal (Elton/Mark D.)
      • BCal is on and working.
      • Need to re-evaluate what value to use for the integration window.
      • Need to take pulser data.
    • FCal (Adesh)
      • 19 bases were replaced on Thursday 26th.
    • TOF (Mark Ito/Beni/Alex. Ostrovidov)
      • Need to find the optimal HV. The HV-scan data from Fall 16 is not as good as first though, in particular because some have been accidentally lost.
      • There was a broken ADC in slot 5. It is now replaced but it has the wrong firmware.
  • Online monitoring (David L.)
    • Added a window with reference plots to the monitoring.
  • Calibration/reconstruction/offline monitoring (Paul Mattione/ Alexander Austregesilo)
    • Software was updated. Alexander will check if the new codes are fine.
  • Level 3 Trigger (David L.)
  • Safety documents (Beni):
    • ESAD was updated to include safety procedure for TRD and is now ready. <strike>
    • RSAD
  • Hall Status/safety (Tom/Tim)
  • Other topics?